OpenStudy (anonymous):

Find the d/dx of the integral of e^t^3 dt from 3 to x^2

OpenStudy (accessdenied):

What have you tried with this problem so far? Any ideas on what we might use?

OpenStudy (anonymous):

I tried the FFT, but I am having trouble solving it because of the log.

OpenStudy (accessdenied):

First fundamental theorem? That sounds like a good idea. And if you could, it'd be great to see the work up to where you got stuck!

OpenStudy (anonymous):

F'(x^2)-F'(3) This is where I am stuck....2xe(^-x^3)-(e(^-3^2))

OpenStudy (primeralph):

Why you stuck?

OpenStudy (accessdenied):

That setup looks good. And then you take the derivative of F(x^2) - F(3). d/dx ( F(x^2) - F(3) ) We take the derivative by terms first. d/dx( F(x^2) - d/dx( F(3) ) But notice that the second term is just a constant. We plug in a value for the variable, so F(3) is just some number. What is the derivative of a constant?

OpenStudy (anonymous):

The derivative of a constant is zero, so d/dx(Fx^2)-d/dx(F(3))=2x-0?

OpenStudy (accessdenied):

We're getting closer. We are left with d/dx( F(x^2) ) after making that second part 0. And I agree that we need chain rule. But I think we just need to be a bit more careful. d/dx(x^2) * f(x^2) This time f(x^2) is from the integrand, we plug in x^2 directly for t here: \(f(t) = e^{t^3} \)

OpenStudy (anonymous):

So f(x^2)=e(^x^2)^3) or e^x^6

OpenStudy (accessdenied):

Yes! We end up with 2x * e^(x^6) now, making sure our chain rule is accounted for. :)

OpenStudy (anonymous):

Thank you so much. I was really over-thinking this problem.

OpenStudy (accessdenied):

Glad to help! :)